The prior probabilities for events A_1 and A_2 are
Answer #1

4)-0 а)Yes. Aand A2 are mutually exclusive since P(A rpna.P@n 4)=P(B/4)P(4)=0.20x 040-0.08 P(A) 10.442>MBA A2)-RB/4)P(4)=0.05